OpenStudy (anonymous):

Solve x + 3y = 9 3x – 3y = –13 a. (–1, ten thirds) b. (1, negative ten thirds) c. (–1, 3 over 10) d. (1, negative 3 over 10)

OpenStudy (callisto):

x + 3y = 9 -(1) 3x – 3y = –13 -(2) (1)+(2) (x + 3y) + (3x – 3y) = 9 + (-13) <- Can you solve it?

OpenStudy (anonymous):

@Callisto -3(x + 3y = 9) 3x – 3y = –13 -3x - 9y = -27 3x – 3y = –13 -12y = -58 y = 58 --- 12

OpenStudy (callisto):

It's good that you did it using your way~ Though, there are some problems.. -3x - 9y = -27 <- correct 3x – 3y = –13 <-correct -12y = -58 <- something wrong with the right side, can you check what -27-13 is? y = 58 <- incorrect --- 12

OpenStudy (anonymous):

-12y = 14

OpenStudy (callisto):

No... what is -27-13??

OpenStudy (anonymous):

-40

OpenStudy (callisto):

Yes. So, for the right side, you get -40, and for the left, you got -12y. Can you solve it? (-12y=-40)

OpenStudy (anonymous):

-40 -20 -10 --- = ------ = ----- -12 - 6 -3

OpenStudy (callisto):

-10/-3 can be further simplified by cancelling the common factor -1. What is the final answer you got for y?

OpenStudy (anonymous):

y = 10 - ___ 3 instead of replacing Y in the other equations and such, since there's only one answer choice with this number, can't we determine that it is the correct answer?

OpenStudy (callisto):

No negative sign indeed! negative/ negative = positive

OpenStudy (anonymous):

ohhh, okay, so the answer is A. ?